# Motion blur; Blind deblurring; Blind motion deblurring; Deep learning

### v1

@@ 2.1.2 End-to-end CNN-based Blind Motion Deblurring Methods
2.2 RNN-based Blind Motion Deblurring Methods
Recurrent Neural Networks (RNNs) are commonly used for handling sequential data such as text, language, and time series, demonstrating exceptional performance when dealing with data that exhibit temporal or sequential patterns Liping et al. (2022). Due to their memory-based neural network architecture, RNNs have found applications in image processing as well. For instance, Zhang et al. Zhang et al. (2018) proposed an end-to-end spatially variant RNN network for deblurring in dynamic scenes. As illustrated in Figure 6, the weights of RNNs are learned by deep CNNs. They analyzed the relationship between spatially variant RNNs and the deconvolution process, demonstrating the ability of spatially variant RNNs to simulate the deblurring process. Models trained using the proposed RNNs are notably smaller and faster. To better utilize the properties of spatially varying RNNs, Ren et al. Ren et al. (2021) extend a unidirectionally connected one-dimensional spatially varying RNN to a two-dimensional spatial RNN with three-way connectivity. 2D RNNs can capture a larger receptive field and learn denser propagation between neighboring pixels compared to 1D RNNs, further improving deblurring performance. Tao et al. Tao et al. (2018) proposed a multi-scale deep neural network for single image deblurring based on convolutional long short-term memory (Conv LSTM), effectively utilizing previous frames and image features. Dongwon Park et al. Park et al. (2020) developed a multi-temporal recurrent neural network (MT-RNN) based on recursive feature mapping for blind single-image deblurring, progressively deblurring a single image in an iterative process. In practical applications, CNNs and other models are often combined with RNNs to compensate for certain limitations in image processing, such as sequence length restrictions and computational complexity.
@@ 2.2 RNN-based Blind Motion Deblurring Methods
2.3 GAN-based Blind Motion Deblurring Methods
With GANs Goodfellow et al. (2014) showing promising results in computer vision tasks, they have also been applied to image deblurring tasks. The basic process of image deblurring based on GANs is illustrated in Figure 7. The generator learns to restore clear images from blurry ones, while the discriminator assesses whether the generated clear images are real, providing feedback to adjust the training of both the generator and discriminator.
@@ 2.3 GAN-based Blind Motion Deblurring Methods
Ramakrishnan et al. Ramakrishnan et al. (2017) proposed a depth filter based on GAN. For batter restoration performance, it combines global jump connections and dense structure. Additionally, the model omits the blur kernel estimation process, significantly reducing the testing time required for practical applications. Lin et al. Lin et al. (2019) proposed an adversarial blurred region mining and refinement method. This method can drive CGAN Mirza and Osindero (2014) to learn real sharp image distributions globally and locally for blind image deblurring. Based on GANs, O.Kupyn et al. Kupyn et al. (2018) proposed an end-to-end image deblurring network termed deblurGAN.DeblurGAN was trained based on WGAN Arjovsky et al. (2017) with gradient penalty Gulrajani et al. (2017) and perceptual loss Johnson et al. (2016), achieving notable deblurring effects. Based on DeblurGAN, O. Kupyn et al. Kupyn et al. (2019) proposed DeblurGAN-v2. This model incorporates a feature pyramid network within the generator. The discriminator employs a relativistic discriminator and uses least squares loss to assess the generated images on both global and local scales. Moreover, recognizing the impact of backbone network selection on deblurring quality and efficiency, DeblurGAN-v2 utilizes an embedded architecture that allows the use of different backbone networks for feature propagation. Compared to DeblurGAN, DeblurGAN-v2 demonstrates significant improvements in both performance and efficiency. Peng et al. Peng et al. (2023) also proposed the MobileNetDeblur-GAN (MND-GAN) deblurring algorithm based on the DeblurGAN algorithm. This algorithm utilizes MobileNet as a feature extractor to build a feature pyramid network. Additionally, it employs a multi-scale discriminator to jointly train the generator, reducing the computational cost of the generator while ensuring deblurring quality.
@@
To further enhance image deblurring performance, Zhang et al. Zheng et al. (2019) proposed a multi-scale deblurring network based on GANs. The model first restores the edge information of the blurred image and then utilizes clear edges to guide a multi-scale network for image deblurring. Starting from the blurry images themselves, Zhang et al. Zhang et al. (2020) proposed an image deblurring approach that combines two GAN models: a BGAN subnetwork for learning the image blurring process to generate motion-blurred images and a DBGAN for learning the restoration of motion-blurred images. By learning how to blur images, this method can better remove blurriness.
@@
Supervised learning relies on a large amount of training data, but in reality, there are few real blurred-clear image pairs available for collection. Madam et al. Nimisha et al. (2018) proposed an unsupervised motion image deblurring method based on GANs. The model employs GANs to learn the mapping relationship from blurred images to clear images. To prevent mode collapse and maintain consistency between deblurred and clear image gradients, the model incorporates a deblurring CNN module and a gradient module to constrain the solution space and guide the deblurring process. GAN-based methods may encounter issues during training such as model collapse and gradient disappearanceMescheder et al. (2018), potentially resulting in artifacts and loss of details in the reconstructed images. Building upon the cycle-consistent generative adversarial network (CycleGAN) Zhu et al. (2017), Wen et al. Wen et al. (2021) proposed an unsupervised image deblurring method using a multi-adversarial optimization CycleGAN to address artifacts in high-resolution image generation. This method iteratively generates clear high-resolution images through multiple adversarial constraints. Considering the significant role of edge restoration in the structure of deblurring blurry images, this method introduces a structural-aware strategy. It utilizes edge information to guide the network in deblurring, enhancing the network’s ability to recover image details. Zhao et al. Zhao et al. (2023) introduced a contrast-guided multi-attention CycleGAN unsupervised algorithm for motion deblurring. It improved the deblurring performance of unsupervised algorithms. Lu et al. Lu et al. (2019) proposed an unsupervised method for single image deblurring without pairing training images termed UID-GAN. The method introduces a disentangled framework to split the content and blur features of blurred images, which improves the deblurring performance.
Due to the current drawbacks of unsupervised learning, such as large model sizes, lengthy inference times, and strict constraints on image resolution and domains, Zhao et al. Zhao et al. (2022) proposed an unsupervised blind image deblurring model termed frequency-domain contrastive loss constrained lightweight CycleGAN (FCL-GAN). It designed lightweight domain transformation units and parameter-free frequency domain contrast units to ensure the algorithm’s "lightweight" and "real-time" operation, overcoming the limitations of the generally large models and complex inference in unsupervised methods. Additionally, this model enhanced deblurring performance by introducing frequency-domain contrastive learning. Considering the practical application of image deblurring algorithms in real-world scenarios, Liu et al. Liu et al. (2022) proposed a lightweight generative adversarial network termed Ghost-DeblurGAN. While this algorithm may not reach the advanced levels of deblurring, its advantage lies in its small model size and short inference time.
@@ 2.4 Transformer-based Blind Motion Deblurring Methods
The Transformer, renowned for its powerful attention mechanism, utilizes self-attention (SA) to capture long-range dependencies among patches in image sequences and adaptability to given input content Vaswani et al. (2017); Liu et al. (2021); Zhou et al. (2022); Raghu et al. (2021). When applied in image deblurring, it can achieve significant improvements in deblurring effectiveness.
Wang et al. Wang et al. (2022) proposed an efficient and effective Transformer-based image restoration architecture called U-former by embedding local enhancement window Transformer blocks into a U-shaped structure Ronneberger et al. (2015). They inserted learnable multi-scale restoration modulators into their decoder to enhance the quality of image restoration. The basic architecture is depicted in Figure 8. Although Uformer has shown promising performance in image restoration tasks, the heavy computation involved in self-attention restricts the application of the transformer within the encoding-decoding framework to limited depths. This inevitably affects the effectiveness of image restoration. Building upon U-former, Feng et al. Ji et al. (2021) proposed a deep encoder-decoder image restoration network called U2-former, with Transformer as the core operation. This model employs two nested U-shaped structures to facilitate feature mapping at different scales, resulting in an enhanced image deblurring effect. To optimize computational efficiency, the model utilizes a feature filtering mechanism to discard low-quality features, thereby reducing unnecessary computations.
@@
To effectively eliminate non-uniform blur, Tsai et al. Tsai et al. (2022) proposed an efficient Transformer-based image deblurring algorithm termed Stripformer. This method reweights image features by computing intra-band and inter-band attention to address blur of different sizes and orientations. It achieves notable deblurring results without requiring extensive training data, enhancing the removal of non-uniform blur. Zamir et al. Zamir et al. (2022) proposed a Transformer-based image restoration network termed Restormer. This method employs a multi-head attention mechanism for feature fusion to handle high-resolution images in image restoration tasks. Additionally, this method utilizes multi-scale local-global representation learning to effectively utilize distant contextual information. It simultaneously reduces computations while enhancing the effectiveness of image restoration. Kong et al. Kong et al. (2023) proposed an asymmetric network based on an encoder-decoder architecture. They designed an effective frequency-domain self-attention solver in the decoder, reducing spatial complexity and computational demands through element-wise operations. Additionally, they employed a discriminative frequency-based feedforward network (DFFN) to discern which low-frequency and high-frequency feature information should be retained for potential clear image restoration. Recently, Zhong et al. Zhong et al. (2023) proposed a Blur Interpolation Transformer (BiT) to reveal potential temporal correlations in blurry encodings. Based on multi-scale residual Swin Transformer blocks, they introduced dual-ended temporal supervision and a time-symmetric integration strategy to generate effective time-varying motion-rendered features, effectively generating clear image sequences from blurry inputs. Zou et al. Zou and Ma (2023) explored the significant roles of edges and textures in image deblurring and proposed a method called Edgeformer. This method utilizes edge information to enhance the correlation between pixels. By employing an edge-enhanced feed-forward network, it distinguishes between edge and texture features, resulting in improved image deblurring. To fully leverage both local and non-local features in images, Wu et al. Wu et al. (2023) introduced the hierarchical pyramid attention transformer (HPAT) network, based on a transformer-based multi-level network structure. It employs hierarchical attention fusion modules (HAFM) to extract global contextual information, enhancing the feature representation capability. Using a fusion feed-forward neural network (F3N) for feature fusion has improved the deblurring effectiveness. Liang et al. Liang et al. (2023) used deep features extracted from pre-trained visual transformers (ViT) Dosovitskiy et al. (2020) to encourage the recovered image to become clear without sacrificing quantitative performance. They also proposed local MAE perceptual loss and global distribution perceptual loss to guide image deblurring.
@@
The existing image deblurring methods mainly focus on global deblurring, which may affect the clarity of locally blurred images. Li et al. Li et al. (2023) proposed a local dynamic motion deblurring transformation network (LDM-ViT) based on adaptive window pruning transformer blocks (AdaWPT). In order to focus the network attention on localized blur regions, AdaWPT removes unnecessary windows. Only the blurred regions are allowed to participate in the deblurring process, reducing the amount of unnecessary computation. Additionally, using labeled local blur image datasets to train the predictor can be able to remove local motion blur effectively. Yan et al. Yan et al. (2023) introduced a Transformer-based dynamic scene image deblurring network called SharpFormer. They designed a Local-preserving Transformer block (LTransformer) that learns local features while retaining global representations. To effectively apply LTransformer blocks to mid-resolution features, they used dynamic convolution blocks to adaptively handle input non-uniform blur, enhancing the quality of deblurred images.

### v2

@@ 2.2 RNN-based Blind Motion Deblurring Methods
2.3 GAN-based Blind Motion Deblurring Methods
With GANs [62] showing promising results in computer vision tasks, they have also been applied to image deblurring tasks. The basic process of image deblurring based on GANs is illustrated in Figure 7. The generator learns to restore clear images from blurry ones, while the discriminator assesses whether the generated clear images are real, providing feedback to adjust the training of both the generator and discriminator. GAN-based methods can be divided into two categories, supervised and unsupervised models, based on different training methods and data dependencies.
@@ 2.3 GAN-based Blind Motion Deblurring Methods
Supervised methods: Supervised methods use blur-sharp image pairs as supervisory signals for neural network training to guide model updating and learning. Supervised motion deblurring can more accurately learn the nonlinear mapping between blurred and sharp images, better removing blur and restoring details. Ramakrishnan et al. [96] proposed a depth filter based on GAN. For batter restoration performance, it combines global jump connections and dense structure. Additionally, the model omits the blur kernel estimation process, significantly reducing the testing time required for practical applications. Lin et al. [97] proposed an adversarial blurred region mining and refinement method. This method can drive CGAN [98] to learn real sharp image distributions globally and locally for blind image deblurring. Based on GANs, O.Kupyn et al. [99] proposed an end-to-end image deblurring network termed deblurGAN. DeblurGAN was trained based on WGAN [100] with gradient penalty [101] and perceptual loss [102], achieving notable deblurring effects. Based on DeblurGAN, O. Kupyn et al. [103] proposed DeblurGAN-v2. This model incorporates a feature pyramid network within the generator. The discriminator employs a relativistic discriminator and uses least squares loss to assess the generated images on both global and local scales. Moreover, recognizing the impact of backbone network selection on deblurring quality and efficiency, DeblurGAN-v2 utilizes an embedded architecture that allows the use of different backbone networks for feature propagation. Compared to DeblurGAN, DeblurGAN-v2 demonstrates significant improvements in both performance and efficiency. Peng et al. [104] also proposed the MobileNetDeblur-GAN (MND-GAN) deblurring algorithm based on the DeblurGAN algorithm. This algorithm utilizes MobileNet as a feature extractor to build a feature pyramid network. Additionally, it employs a multi-scale discriminator to jointly train the generator, reducing the computational cost of the generator while ensuring deblurring quality.
@@
To further enhance image deblurring performance, Zhang et al. [105] proposed a multi-scale deblurring network based on GANs. The model first restores the edge information of the blurred image and then utilizes clear edges to guide a multi-scale network for image deblurring. Starting from the blurry images themselves, Zhang et al. [106] proposed an image deblurring approach that combines two GAN models: a BGAN subnetwork for learning the image blurring process to generate motion-blurred images and a DBGAN for learning the restoration of motion-blurred images. By learning how to blur images, this method can better remove blurriness. Pan et al. [107] applied the physical model to generative adversarial networks for image restoration, and proposed a GAN model constrained by the physical model. The physical model is derived from the image formation process of the underlying problem. Considering the practical application of image deblurring algorithms in real-world scenarios, Liu et al. [108] proposed a lightweight generative adversarial network termed Ghost-DeblurGAN. While this algorithm may not reach the advanced levels of deblurring, its advantage lies in its small model size and short inference time.
@@
Unsupervised methods: Supervised learning relies on a large amount of training data, but in reality, there are few real blurred-clear image pairs available for collection. Madam et al. [109] proposed an unsupervised motion image deblurring method based on GANs. The model employs GANs to learn the mapping relationship from blurred images to clear images. To prevent mode collapse and maintain consistency between deblurred and clear image gradients, the model incorporates a deblurring CNN module and a gradient module to constrain the solution space and guide the deblurring process. GAN-based methods may encounter issues during training such as model collapse and gradient disappearance[110], potentially resulting in artifacts and loss of details in the reconstructed images. Building upon the cycle-consistent generative adversarial network (CycleGAN) [111], Wen et al. [112] proposed an unsupervised image deblurring method using a multi-adversarial optimization CycleGAN to address artifacts in high-resolution image generation. This method iteratively generates clear high-resolution images through multiple adversarial constraints. Considering the significant role of edge restoration in the structure of deblurring blurry images, this method introduces a structural-aware strategy. It utilizes edge information to guide the network in deblurring, enhancing the network’s ability to recover image details. Zhao et al. [113] proposed a contrast-guided multi-attention CycleGAN unsupervised algorithm for motion deblurring. It improved the deblurring performance of unsupervised algorithms. Lu et al. [114] proposed an unsupervised method for single image deblurring without pairing training images termed UID-GAN. The method introduces a disentangled framework to split the content and blur features of blurred images, which improves the deblurring performance. Due to the current drawbacks of unsupervised learning, such as large model sizes, lengthy inference times, and strict constraints on image resolution and domains, Zhao et al. [115] proposed a lightweight CycleGAN (FCL-GAN) with frequency domain contrast loss constraints. FCL-GAN designed lightweight domain transformation units and parameter-free frequency domain contrast units to ensure the algorithm’s "lightweight" and "real-time" operation, overcoming the limitations of the generally large models and complex inference in unsupervised methods.
@@ 2.4 Transformer-based Blind Motion Deblurring Methods
In recent years, models based on transformers have demonstrated tremendous potential in visual tasks [116, 117]. The Transformer, renowned for its powerful attention mechanism, utilizes self-attention (SA) to capture long-range dependencies among patches in image sequences and adaptability to given input content [118, 119, 120]. When applied in image deblurring, it can achieve significant improvements in deblurring effectiveness. The Transformer model itself possesses the capability to handle information of varying scales and ranges, allowing it to flexibly adapt to different types of blur problems. According to the specific image blurring situation and deblurring requirements the transformer-based deblurring methods are divided into two categories: global and local deblurring.
Global deblurring methods: Global deblurring methods are mostly applied to the blurring of the whole image produced by camera shake. Global deblurring methods do not distinguish between sharp and blurred regions and consider the whole image as a blurred whole. The same deblurring process is done for blurred and sharp regions. Wang et al. [121] proposed an efficient and effective Transformer-based image restoration architecture called U-former by embedding local enhancement window Transformer blocks into a U-shaped structure [122]. They inserted learnable multi-scale restoration modulators into their decoder to enhance the quality of image restoration. The basic architecture is depicted in Figure 8. Although Uformer has shown promising performance in image restoration tasks, the heavy computation involved in self-attention restricts the application of the transformer within the encoding-decoding framework to limited depths. This inevitably affects the effectiveness of image restoration. Building upon U-former, Feng et al. [123] proposed a deep encoder-decoder image restoration network called U2-former, with Transformer as the core operation. This model employs two nested U-shaped structures to facilitate feature mapping at different scales, resulting in an enhanced image deblurring effect. To optimize computational efficiency, the model utilizes a feature filtering mechanism to discard low-quality features, thereby reducing unnecessary computations.
@@
To effectively eliminate non-uniform blur, Tsai et al. [124] proposed an efficient Transformer-based image deblurring algorithm termed Stripformer. This method reweights image features by computing intra-band and inter-band attention to address blur of different sizes and orientations. It achieves notable deblurring results without requiring extensive training data, enhancing the removal of non-uniform blur. Zamir et al. [125] proposed a Transformer-based image restoration network termed Restormer. Restormer employs a multi-head attention mechanism for feature fusion to handle high-resolution images in image restoration tasks. Additionally, this method utilizes multi-scale local-global representation learning to effectively utilize distant contextual information. It simultaneously reduces computations while enhancing the effectiveness of image restoration. Kong et al. [126] designed an effective frequency-domain self-attention solver in the decoder, reducing spatial complexity and computational demands through element-wise operations. Additionally, they employed a discriminative frequency-based feedforward network (DFFN) to discern which low-frequency and high-frequency feature information should be retained for potential sharp image restoration. Recently, Zhong et al. [127] proposed a Blur Interpolation Transformer (BiT) to reveal potential temporal correlations in blurry encodings. Based on multi-scale residual Swin Transformer blocks, they introduced dual-ended temporal supervision and a time-symmetric integration strategy to generate effective time-varying motion-rendered features, effectively generating clear image sequences from blurry inputs. Zou et al. [128] explored the significant roles of edges and textures in image deblurring and proposed a method called Edgeformer. This method utilizes edge information to enhance the correlation between pixels. By employing an edge-enhanced feed-forward network, it distinguishes between edge and texture features, resulting in improved image deblurring. To fully leverage both local and non-local features in images, Wu et al. [129] introduced the hierarchical pyramid attention transformer (HPAT) network, based on a transformer-based multi-level network structure. It employs hierarchical attention fusion modules (HAFM) to extract global contextual information, enhancing the feature representation capability. Using a fusion feed-forward neural network (F3N) for feature fusion has improved the deblurring effectiveness. Liang et al. [130] used deep features extracted from pre-trained visual transformers (ViT) [131] to encourage the recovered image to become clear without sacrificing quantitative performance. They also proposed local MAE perceptual loss and global distribution perceptual loss to guide image deblurring.
@@
Local deblurring methods: The existing image deblurring methods mainly focus on global deblurring, which may affect the clarity of locally blurred images. The local deblurring method mainly focuses on the situation where the foreground and background are moving relative to each other. The foreground is blurred while the background is sharp. Local deblurring methods will separate the blurred foreground region by taking mask and other operations to ensure that there are no adverse effects on the sharp regions during the deblurring process. Li et al. [132] proposed a local dynamic motion deblurring transformation network (LDM-ViT) based on adaptive window pruning transformer blocks (AdaWPT). In order to focus the network attention on localized blur regions, AdaWPT removes unnecessary windows. Only the blurred regions are allowed to participate in the deblurring process, reducing the amount of unnecessary computation. Additionally, using labeled local blur image datasets to train the predictor can be able to remove local motion blur effectively. Yan et al. [133] proposed a Transformer based deblurring network for dynamic scene images termed SharpFormer. To effectively model the detailed features of the image, the model uses locality preserving Transformer (LTransformer) blocks to learn the local features. In addition, the model uses dynamic convolutional blocks to adaptively deal with the non-uniform blurring of the input images, improving the quality of the deblurred image.
2.5 Advantages and limitations of the four architectures.
In image deblurring, the four architectures CNN, RNN, GAN, and Transformer have their own advantages and limitations as shown in Table LABEL:tab:advantages :
CNN is widely used in image processing to capture local features and spatial information [134, 135]. When trained on large datasets, deblurring methods based on CNN perform high efficiency and generalization ability. CNN demonstrates adaptability to different types of blurry scenarios, enabling it to handle various forms of blur such as motion blur and defocus blur. The CNN structure is simple and suitable for image denoising and deblurring tasks [136]. However, CNN performance typically relies on extensive training data, requiring a substantial number of clear and blurry image pairs for effective training. Additionally, CNN may be constrained by a fixed-size receptive field, potentially resulting in suboptimal performance in image deblurring tasks involving global information or long-range dependencies. In order to solve the problem of limited receptive field, the more widely used method is to use the dilated convolution.
RNN can effectively capture temporal information in images, aiding in handling blurs with time dependencies, such as motion blur. However, there may be gradient vanishing or explosion problems when dealing with long-term dependencies [137]. Moreover, for image deblurring tasks, RNN is not well suited to capture spatial information [95]. Therefore, RNNs are generally combined with other structures to accomplish image deblurring tasks.
GAN enhances the realism of image generation through adversarial training, providing a significant advantage in deblurring tasks. The GAN generator can produce diverse outputs, helping to deal with different scenarios and different types of blurring, thereby improving the model’s generalization ability. In contrast to CNN and Transformer methods, the generator training of GAN does not depend entirely on data samples, and its gradient update information comes from the discriminator. GAN can learn the task of image deblurring through unsupervised learning from a large amount of unlabeled data. Image motion deblurring methods based on GAN improve the accuracy of parameter fitting based on the principles of game theory. However, the training process may be unstable, requiring a balance between generator and discriminator training [138]. There may also be issues such as pattern crashes or non-convergence of training patterns.
Transformer is able to capture global information and solve the remote spatial dependence problem [133], and has advantages in processing for some image tasks that require long-distance dependence. Transformer can capture the global information of input images, aiding in understanding the overall blurry structure and image content for better restoration of clear images. The introduction of positional encoding in Transformer enables effective handling of spatial relationships in images, making the model more targeted in dealing with blur. The self-attention mechanism of Transformer allows it to dynamically adjust its focus on different regions of the image, helping to address varying degrees of blur. However, the image deblurring task involves the processing of a large number of pixels and has a high computational cost. Compared to traditional CNN structures, Transformer models typically have a larger parameter count [117], which may result in higher training and inference costs.
Methods Pub. Category Overview Advantages Limitations
DeepDeblur [79] CVPR 2017 CNN Removing complex motion blur from coarse to fine using a multi-scale recurrent network End-to-end removal of complex motion blur without the need for kernel estimation The mode has a large number of parameters,making training difficult
DMPHN [87] CVPR 2019 CNN Estimated by segmenting the image into multiple blocks and then removing blur at different scales Overcoming the limitations of depth deblurring models on superposition depths Feature segmentation may lead to discontinuous contextual information
PSS-NSC [139] CVPR 2019 CNN Multi-scale CNN deblurring networks are analyzed and a parameter-selective sharing scheme is proposed Selective parameter sharing can improve the learning efficiency and generalization of the model The introduction of new network structures and hopping connection mechanisms may have increased computational costs
DGN [140] TIP 2020 CNN An effective depth-guided network consisting of an image deblurring branch and a depth refinement branch is proposed The use of depth maps allows for a better understanding of scene changes and facilitates the use of depth information Deblurring quality depends on the accuracy and quality of the depth map
MSCAN [141] TSCVT 2020 CNN A multi-scale channel attention network is proposed for more robust feature representation and higher quality image recovery Further taking advantage of multi-scale models, codec modules and residual block structures The model has many parameters and is difficult to train
SDWNet [142] ICCV 2021 CNN A network structure combining extended convolution and wavelet reconstruction is proposed for image deblurring Dilated convolution results in a larger perceptual field and reduces the loss of texture information during sampling Introduction of wavelet transform increases the complexity of the network
MIMOU-Net [85] ICCV 2021 CNN Introducing asymmetric feature fusion for gradual recovery of clear images by coder-decoder Handling multi-scale blurry with low computational complexity Improvement of deblurring accuracy increases computational effort and time consuming
MPRNet [90] CVPR 2021 CNN Multi-stage architecture, injecting supervision at each stage to incrementally improve degraded inputs Horizontal connections exist between feature processing blocks to avoid information loss High dependence on computing power
MSSNet [143] ECCV 2022 CNN A deep learning-based multi-scale stage network from coarse to fine is proposed Multi-scale stage configurations and inter-scale information propagation contribute to the network’s deblurring performance Hyperparameter adjustment is more complex
HINet [144] CVPRW 2022 CNN Propose semi-instantiation to one piece using multi-stage architecture for image recovery task HINet architecture generalizes well and has short inference time Requires a large number of image pairs for training
BANet [145] TIP 2022 CNN Localization of blur direction and size by blurry perception for adaptive blur removal Short inference time can well support subsequent real-time applications Large amounts of diversity data are needed for training to adapt to dynamic scene blurring
IRNeXt [146] ICML 2023 CNN Integration of multi-stage mechanisms into a u-shaped network to remove different sizes of blur in a coarse-to-fine manner The local attention module improves restoration quality by weighting high-frequency information The performance improvement of deblurring depends on the number of MSM branches
MRLPFNet [84] ICCV 2023 CNN A simple and effective multi-scale residual low-pass filtering network Effectively model both high-frequency and low-frequency information It moderately increases the running time
SVRNN [64] CVPR 2018 RNN End-to-end trainable spatially-variant RNN for dynamic scene deblurring The pixel-by-pixel weights of the RNN are learned by the deep CNN, which helps to remove blurring Need for a larger feeling field, dealing with large areas and spatial variations
SRN [81] CVPR 2018 RNN Blur removal is achieved using a decoder with skip connections and parameter sharing across three scale Sharing network weights at different scales, thus significantly reducing training complexity Ignoring the variations between feature scales can easily saturate the network’s performance
MT-RNN [91] ECCV 2019 RNN Gradual removal of non-uniform blurring over multiple iterations using a self-looping MT structure Contextual information is continuous, which facilitates the handling of blurring in dynamic scenes Inflexible progressive training may not remove non-uniform blurring very well
Optical Flow network [29] TCSVT 2022 RNN Blur removal using optical flow guided spatially variable RNNs Optical flow can provide effective information for deblurring dynamic scenes Inaccurate optical flow will make it more difficult to recover a clear image
DeblurGAN [99] CVPR 2018 GAN Conditional GAN-based network for generating realistic deblurred images Uniform and non-uniform motion blur images with less blurring can be recovered Neglecting the nonlinearity characteristics of dynamic scenes blur, the network sensory field is small and homogeneous
DeblurGAN-V2 [103] ICCV 2019 GAN Use feature pyramid networks and extensive backbone networks for better accuracy Improved image restoration efficiency and enhanced real-time performance Single network structure, which tends to lead to saturation of the network model
DBGAN [106] CVPR 2020 GAN Better deblurring by combining learning to generate blurred images and deblurring images Synthetic blurred images more accurately mimic real blur for better deblurring Network training is complex
Physics-Based GAN [107] TPAMI 2021 GAN Applied the physical model to generative adversarial networks for image restoration The physical model is derived from the image formation process and facilitates image recovery Less effective for complex low quality images
CycleGAN [111] ICCV 2021 GAN Blind motion deblurring based on multi-adversarial mechanism for iterative generation of high-resolution images Add multi-scale edge constraint function to enhance network deblurring and detail preservation capability Presence of large color differences, severe artifacts and model redundancy
FCL-GAN [115] MM 2022 GAN Constraining lightweight network CycleGAN driven tasks by frequency domain contrast losses Better deblurring of images in a lightweight and real-time manner Deblurring effect does not reach the supervised model SOTA level
Ghost-DeblurGAN [108] IROS 2022 GAN Introducing GhostNet for designing lightweight deblurring networks Suitable for real-time image deblurring, small model size, fast inference time Deblurring effect fails to reach SOTA level
Uformer [121] CVPR 2021 Transformer Building hierarchical codec networks with the Transformer module Ability to capture local and global dependencies Self-attentive layers can only be applied to a limited depth
Restormer [125] CVPR 2022 Transformer Codecs for learning multi-scale local-global representations of high-resolution images Ability to model global connectivity with high computational efficiency Image recovery does not fully utilize spatial information
Stripformer [124] ECCV 2022 Transformer Re-weighting image features horizontally and vertically to capture blurred features in different directions Low memory footprint and computational cost, does not rely on large amounts of training data Scaled dot product attention often requires complex matrix multiplication
FFTformer [126] CVPR 2023 Transformer Combining a self-attentive solver with a feed-forward neural network based on discriminative frequency domain to remove blur Lower space and time complexity and more effective and efficient in image deblurring May not generalize well to different types of images and blurs
BiT [127] CVPR 2023 Transformer Introducing Double-Ended Temporal Supervision and Temporal Symmetry Integration Strategies to Generate Effective Time-Varying Motion Rendering Features Capable of effectively handling blurring at different scales and fusing information from neighboring frames Limited discrete supervision does not respond well to different realities
Sharpformer [133] TIP 2023 Transformer Adaptive handling of non-uniform blurring with dynamic convolutional blocks by fully utilizing global-local features Able to consistently improve the quality of deblurring due to CNN-based deblurring methods alone Higher computational complexity
3 Datasets and Evaluation Metrics
3.1 Datasets
Most neural network-based image deblurring methods typically require paired images for training. Existing blurry image datasets can be roughly divided into two categories: synthetic datasets and real datasets. Synthetic datasets usually generate blurred images by frame averaging. While Real datasets capture blurred image using a long exposure time. Synthetic datasets include the Köhler dataset, Blur-DVS dataset, GoPro dataset, and HIDE dataset, among others. Real image datasets comprise the RealBlur dataset, RsBlur dataset, ReLoBlur dataset, and so on. As shown in Figure 10, these are example images from some commonly used datasets. Table 2 summarizes and outlines the basic information of select datasets. We additionally provide bar graphs to compare the size of the datasets, as shown in Figure 9.
Köhler et al. Dataset: To simulate real camera shake, Köhler et al. [147] used a Stewart platform with six degrees of freedom to replay the camera’s motion trajectory. During the replay of the motion trajectory, a series of clear images and long-exposure blurry images were captured. The dataset comprises 4 clear images and 48 corresponding blurry images for 12 motion trajectories.